OpenStudy (anonymous):

f(x) = sqrt( 1- X^2 ) . f: R -> R find domain of f?

OpenStudy (experimentx):

sqrt( 1- X^2 ) <---- this term for values of X must be a real number.

OpenStudy (anonymous):

1 - \[x^{2}\] must be \[\ge\] 0

OpenStudy (experimentx):

yes.

OpenStudy (anonymous):

omg.

OpenStudy (anonymous):

experiment plz solve

OpenStudy (anonymous):

\[1 - x^{2}\] >= 0 find all values for x

OpenStudy (experimentx):

Oops sorry , i thought x^2 - 1,

OpenStudy (anonymous):

try again San..... :)

OpenStudy (experimentx):

yeah ... i didn't see that equation properly. 1 - x^2 >=0 1 >= x^2 so x^2 must always be less or equal to one. it can have that value in betwen -1 and 1, so this is your domain. it must be x<= 1 {x:x<=1 or x >= -1} or [-1, 1]

OpenStudy (anonymous):

ohh.. i got your answer..

OpenStudy (anonymous):

but one thing more...

OpenStudy (anonymous):

x^2 <= 1 what if we sqrt both side ?

OpenStudy (experimentx):

??

OpenStudy (anonymous):

x^2 <= 1 what if we sqrt both side ?

OpenStudy (experimentx):

you will get x<= -1 and x <= +1 but -3 will not work on our condition, will it??

OpenStudy (experimentx):

the best way to do this would be x^2 <= 1 or, |x^2| <= 1 or, -1 <= |x^2| <= 1 or, -1 <= x <= 1

OpenStudy (anonymous):

but experiment... we know that if we sqrt x^2 we get | x | m i right ?

OpenStudy (anonymous):

if we take modulus both sides, doesn't it effect the inequality?

OpenStudy (experimentx):

I don't think so. that would imply |x| <= 1 and |x| <= -1

OpenStudy (experimentx):

in this case yes, it will not affect. because |x^2| and x^2 will always be positive no matter what the value of x is.

OpenStudy (experimentx):

unless it's non real.

OpenStudy (anonymous):

ohhhh.. i got u :)

OpenStudy (anonymous):

thanks :) alot i will be back with new problem :) :) thanks brooooooooooooooooooooooooooooooooooooooo

OpenStudy (experimentx):

sure ... you are welcome.